How to Simply Square Roots;


1. 25/64
2. 36
3. 3*6
4. 50/10
5. 12/9
6. 64
7. 40
8. 99
9. 196
10. 32
11. 90
12. 324
13. 24
14. 84

Answer by checkley71(8403)

sqrt(25/36)=5/6
sqrt36=6
sqrt(3*6)=sqrt18=sqrt9*2=3sqrt2
sqrt(50/10)sqrt(25*2)/10=5(sqrt2/10)
sqrt(12/9)=sqrt(4*3/3*3)=2/3sqrt3
I'll leave the rest for to practice with.